Circular No. 8844

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 A4–A7, C/2007 B4–B6, C/2007 C3–C13 (SOHO) The “discovery” data for additional near-sun comets found on SOHO

website images are tabulated below (cf. IAUC 8819) — all being Kreutz sungrazers except C/2007 A6 and C/2007 C10 (Meyer group) and C/2007 A7, C/2007 C7, and C/2007 C12 (no known group). C/2007 A4 and C/2007 A5 were small and stellar in appearance in C3-coronagraph images; in C2 images, C/2007 A4 was tailless and slightly diffuse. C/2007 A6 and C/2007 C10 were small and stellar in appearance, of mag ∼ 7. C/2007 A7 was stellar in appearance, peaking at mag ∼ 6 but dying out rapidly. C/2007 B6 was brighter (mag ∼ 5) with a hint of a tail. C/2007 C3 reached mag ∼ 4 and showed a very short, faint tail. C/2007 C6 reached mag ∼ 3; it showed a very faint, thin tail ∼ 0^{o}.5 long in C2 images. C/2007 C7 peaked at mag 7.5 and faded gradually, being tiny and stellar in appearance. C/2007 C11 was stellar in appearance, reaching mag ∼ 6. C/2007 C12 was stellar in appearance with a hint of a tail; it brightened (reaching mag ∼ 7) then died rapidly. C/2007 C13 was brighter (peaking at mag ∼ 3.5), showing a short, “headless” tail ∼ 20′ long in C2 images. The remaining six objects appeared stellar and very faint with no tail. ‘MU’ = M. Uchina.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 A4 Jan. 4.763 19^{h}10.^{m}9 −25^{o}19′ C3/2 HS 2007-K46
C/2007 A5  9.446 19 31.7 −23 59 C3  HS 2007-K46
C/2007 A6  10.938 19 33.3 −21 21 C2  TH 2007-K53
C/2007 A7  10.314 19 31.0 −23 01 C2  LC 2007-K65
C/2007 B4  18.096 20 13.3 −22 58 C3  HS 2007-K53
C/2007 B5  19.179 20 21.6 −23 05 C3  WX 2007-K53
C/2007 B6  31.113 21 20.2 −20 19 C3/2 HS 2007-K53
C/2007 C3 Feb. 1.138 21 23.8 −20 07 C3/2 MU 2007-K54
C/2007 C4  2.971 21 21.7 −18 14 C3  HS 2007-K54
C/2007 C5  3.679 21 22.5 −17 49 C3  HS 2007-K54
C/2007 C6  5.846 21 44.0 −19 07 C3/2 BZ 2007-K54
C/2007 C7  2.663 21 01.9 −18 33 C2  HS 2007-K65
C/2007 C8  5.821 21 32.4 −17 16 C3  BZ 2007-K65
C/2007 C9  6.779 21 35.0 −16 47 C3  RM 2007-K65
C/2007 C10  7.163 21 26.2 −14 03 C2  HS 2007-K65
C/2007 C11  7.738 21 47.9 −17 17 C3  HS 2007-K66
C/2007 C12  8.404 21 43.2 −17 46 C3  HS 2007-K66
C/2007 C13  14.654 22 22.9 −15 10 C3/2 HS 2007-K66
2007 June 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green